The selling price was $140. If the markup was 20 percent of the purchase price, what was the purchase price and what was the markup (to the nearest penny)?
Answer by jorel555(1290)

Let n be the purchase price. Then:
1.2 x n=140
n=140/1.2=$116.67 purchase price
140-116.67=$23.33 markup amount. ☺☺☺☺